# On the equivalence of all notions of generalized derivations whose domain is a C-algebra

### v1

@@ 2. $(ii)$
We shall next show that $I$ is a closed ideal of $\mathcal{A}$. Namely, take $a\in I$ and $b\in\mathcal{A}$. Clearly, the mapping $FL_{a}L_{b}:\mathcal{A}\rightarrow\mathcal{M}$, $c\mapsto F(abc)$ is continuous, and so $ab=ba\in I$. This means that $I$ is an ideal of $\mathcal{A}$ (recall that $\mathcal{A}$ is commutative). It is well known that $I$ must be self-adjoint [30, Corollary 4.2.10].
Next we show that $I$ is norm-closed. Let $a\in\overline{I}$. Then there exists a sequence $(a_{n})_{n}\subseteq I$ such that $\lim_{n\rightarrow\infty}a_{n}=a$. By the equality $I=J$, in order to show that $a\in I$, it suffices to show that the mapping $L_{a}H:\mathcal{A}\rightarrow\mathcal{M}$ is continuous. It follows from Lemma 2.3$(i)$ and the assumptions that the mapping $L_{a_{n}}H=:\mathcal{A}\rightarrow\mathcal{M}$ is linear for every $c\in\mathcal{A}$. Since $(a_{n})$ is a sequence in $I$, the linear mapping $L_{a_{n}}H:\mathcal{A}\rightarrow\mathcal{M}$ is continuous for all $n\in\mathbb{N}$. It is clear that $\lim_{n\rightarrow\infty}L_{a_{n}}H(c)=\lim_{n\rightarrow\infty}{a_{n}}H(c)=aH(c)=L_{a}H(c),$ for every $c\in\mathcal{A}$, and hence, by the uniform boundedness principle, we obtain that $L_{a}H$ is a continuous linear mapping, and so $a\in J=I$.
@@
Suppose, on the contrary, that ${\mathcal{A}}/{I}$ is an infinite-dimensional C-algebra. Another classical result in C-algebra theory (see [30, Exercise 4.6.13]) proves that in such a case there exists an infinite sequence $(b_{n}+I)_{n}$ of mutually orthogonal non-zero positive elements in ${\mathcal{A}}/{I}$, that is, $(b_{n}+I)(b_{k}+I)=0,$ for all $n\neq k$, $b_{n}\geq 0$ and $b_{n}\neq 0$ for all $n$. By [30, Exercise 4.6.20] we can always lift the sequence $(b_{n}+I)_{n}$ to a sequence $(c_{n})_{n}$ of mutually orthogonal non-zero positive elements in $\mathcal{A}$ satisfying $\pi(c_{n})=b_{n}+I$, for all $n$, where $\pi:\mathcal{A}\rightarrow{\mathcal{A}}/{I}$ is the canonical projection. It follows from the fact that $\pi$ is a -homomorphism that $\pi(c_{n}^{2})=\pi(c_{n})^{2}=(b_{n}+I)^{2}\neq 0$ (observe that $\|(b_{n}+I)^{2}\|=\|b_{n}+I\|^{2}$), and thus $c_{n}^{2}\notin I$ for all $n$. Therefore the mapping $FL_{{c^{2}_{n}}}:\mathcal{A}\rightarrow\mathcal{M}$ must be unbounded. By replacing $c_{n}$ with $\frac{c_{n}}{\|c_{n}\|}$, we can assume that $\|c_{n}\|=1$ for all $n$. By the unboundedness of the mapping $FL_{{c^{2}_{n}}}$, there exists $d_{n}\in\mathcal{A}$ satisfying $\|d_{n}\|\leq 2^{-n},$ and $\|F({{c^{2}_{n}}}d_{n})\|=\|FL_{{c^{2}_{n}}}(d_{n})\|>\|G(c_{n})\|+n$ for all natural $n$ –observe that the sequence $(\|G(c_{n})\|)_{n}$ does not produce any obstacle here–. The elements $c_{n}$ and $d_{n}$ have been chosen to guarantee that the series $\sum_{n\geq 1}c_{n}d_{n}$ is (absolutely) convergent, and its limit $a_{0}=\sum_{n=1}^{\infty}c_{n}d_{n}$ satisfies $\|a_{0}\|\leq 1$ and $c_{m}a_{0}=c_{m}^{2}d_{m}$ for all $m\in\mathbb{N}$. By the hypotheses on $(F,G,H)$ we deduce that
@@
Since $F|_{I}$ is continuous and $\mathcal{A}/I$ is finite dimensional, we deduce from Lemma 2.6 that $F$ is continuous.
@@
The previous theorem generalizes the classical results by Ringrose [38, Theorem 2] and Sakai [43]. The following corollary summarizes some of the conclusions obtained up to now.
@@ 1. $(a)$
    $F$ is a generalized derivation of the first type, i.e., there is a derivation $d:\mathcal{A}\to\mathcal{M}$ such that $F(ab)=F(a)b+ad(b)$ $\forall a,b\in\mathcal{A}.$
@@ 2. $(b)$
    $F$ is a generalized derivation of the second type, i.e., there is an element $\xi\in\mathcal{M}^{**}$ such that $F(ab)=F(a)b+aF(b)-a\xi b$ $\forall a,b\in\mathcal{A}.$
@@ 3. $(c)$
    $F$ is a generalized derivation of the third type, i.e., there are two (non-necessarily linear) maps $G,H:\mathcal{A}\to\mathcal{M}$ such that $F(ab)=G(a)b+aH(b)$ $\forall a,b\in\mathcal{A}.$
Furthermore, if any of the equivalent statements holds the mapping $F$ is automatically continuous.
Let ${Y}$ and ${Z}$ be Banach spaces and let $T:\mathfrak{Y}\rightarrow\mathfrak{Z}$ be a linear mapping. The separating space of $T$ is the set
$\mathfrak{S}(T)=\Big{\{}z\in{Z}\ :\ \exists\ \{y_{n}\}\subseteq{Y}\ such\ that\ y_{n}\rightarrow 0,\hbox{ and }T(y_{n})\rightarrow z\Big{\}}.$
@@
Let $\mathcal{A}$ be a C-algebra and let $\mathcal{M}$ be a Banach $\mathcal{A}$-bimodule. Suppose that $F,G,H:\mathcal{A}\rightarrow\mathcal{M}$ are three mappings, with $F$ linear, such that $(F,G,H)$ behaves like a derivation. The the following statements hold:
@@ 3. $(iii)$
    If $\mathcal{M}=\mathcal{A}$, then $F,G$ and $H$ are bounded linear maps.
@@
$(i)$ According to Theorem 2.8, $F$ is a continuous generalized derivation of the second type, and it follows from Lemma 2.3$(ii)$ that the mappings $G$ and $H$ are linear. We shall show that $G$ and $H$ are continuous. Let $m_{0}\in\mathfrak{S}(H)\subseteq\mathcal{M}$. Then there exists a sequence $\{b_{n}\}\subseteq\mathcal{A}$ such that $\lim_{n\rightarrow\infty}b_{n}=0$ and $\lim_{n\rightarrow\infty}H(b_{n})=m_{0}$. For any arbitrary element $a\in\mathcal{A}$, we have
@@
Let $\mathcal{A}$ be an algebra and let $\mathcal{M}$ be an $\mathcal{A}$-bimodule. We recall that a linear mapping $F:\mathcal{A}\rightarrow\mathcal{M}$ is called an l-generalized derivation if there exists a mapping $d_{l}:\mathcal{A}\rightarrow\mathcal{M}$ (called an associated mapping with $F$) such that $F(ab)=F(a)b+ad_{l}(b)$ for all $a,b\in\mathcal{A}$. Similarly, $F$ is called an r-generalized derivation if there exists a mapping $d_{r}:\mathcal{A}\rightarrow\mathcal{M}$ such that $F(ab)=d_{r}(a)b+aF(b)$ for all $a,b\in\mathcal{A}$. In [20, Example 2.6], we can find an example of an $r$-generalized derivation that is not an $l$-generalized derivation. Clearly, $l$- and $r$-generalized derivations are generalized derivations of the third type.
The next result is therefore a straight consequence of Theorem 2.8 and Corollary 2.10.
@@ 3. Automatic continuity of generalized Jordan derivation of the third type
Appropriate types of generalized Jordan derivations whose domain is a C-algebra are studied in this section. We shall mainly focus on the automatic continuity and the relationships between the different types. Contrary to the conclusions in the previous section, generalized Jordan derivations of the third type constitute a strict subclass of the set of generalized Jordan derivations of the third type (cf. Proposition 3.2 and Remark 3.3). Generalized Jordan derivations of the second type have been considered in [6, 7, 25]. Recall that a linear mapping $F$ from a Banach algebra $\mathcal{A}$ to a Banach $\mathcal{A}$-bimodule $\mathcal{M}$ is a generalized Jordan derivation of the second type if there exists $\xi\in\mathcal{M}^{**}$ such that
@@
where $a\circ b=\frac{1}{2}(ab+ba)$. If $\xi=0$ we find the usual notion of Jordan derivation. Classic results by Ringrose and Johnson assure that every Jordan derivation from a C-algebra to a Banach bimodule is a derivation.
There are some other attempts to defined generalized Jordan derivations in the literature, According to [5], an additive (in this note we shall assume linearity) mapping $T$ from a Banach algebra $\mathcal{A}$ into a left $A$-module $M$ is called a Jordan left derivation if $T(a^{2})=2aT(a)$ for every $a\in\mathcal{A}$. It is shown in [5] that the existence of non-zero Jordan left derivations from a prime ring $\mathcal{R}$ into a 6-torsion free left $\mathcal{R}$-module implies that $R$ is a commutative ring. Vukman introduced in [44] the notion of $(m,n)$-Jordan derivation. Let $m,n$ be two non-negative integers with $m+n\neq 0$. An additive (along this note linear) mapping $T$ from $\mathcal{A}$ into $\mathcal{M}$ is called an $(m,n)$-Jordan derivation if the identity
$(m+n)T(a^{2})=2maT(a)+2nT(a)a$
holds for all $a\in\mathcal{A}$.
Jing and Lu coined for first time the term generalized Jordan derivations in [27]. We shall say that a linear mapping $F$ from a Banach algebra $\mathcal{A}$ to a Banach $\mathcal{A}$-bimodule $\mathcal{M}$ is a generalized Jordan derivation in the sense of Jing and Lu if there exists a (linear) Jordan derivation $\tau:\mathcal{A}\to\mathcal{M}$ satisfying
@@
In this definition the Jordan structure is only considered on the domain algebra. It is worthwhile to mention that if $\mathcal{A}$ and $\mathcal{M}$ are both unital, by taking $b=\mathbf{1}$ in the previous identity we get $F(a)=F(\mathbf{1})a+\tau(a)$ ($a\in\mathcal{A}$), and if $\tau$ is in fact a derivation (as in the case of C-algebras), $F$ is a generalized derivation of the first type.
@@
The following theorem prove the automatic continuity of every generalized Jordan derivation of the third type when the domain is a C-algebra by adapting the proof of Theorem 2.8.
@@
We are now in a position to apply Lemma 2.7 to guarantee that $F|_{I}:{I}\to\mathcal{M}$ is continuous. The continuity of $F$ will follow from Lemma 2.6 if we show that ${\mathcal{A}}/{I}$ is finite-dimensional. Otherwise, as in the proof of Theorem 2.8, there exists an infinite sequence $(b_{n}+I)_{n}$ of mutually orthogonal non-zero positive elements in ${\mathcal{A}}/{I}$, that is, $(b_{n}+I)(b_{k}+I)=0,$ for all $n\neq k$, $b_{n}\geq 0$ and $b_{n}\neq 0$ for all $n$ (see [30, Exercise 4.6.13]). Choose, a sequence $(c_{n})_{n}$ of mutually orthogonal norm-one positive elements in $\mathcal{A}$ satisfying $c_{n}+I=b_{n}+I$, $\|c_{n}^{2}+I\|=\|(b_{n}+I)^{2}\|=\|b_{n}+I\|^{2}\neq 0$, for all $n$ (cf. [30, Exercise 4.6.20]). Clearly, the element $c_{n}^{2}$ does not belong to $I$, and hence the mapping $x\mapsto F(c_{n}^{2}x)$ is unbounded. We can therefore pick $d_{n}\in\mathcal{A}$ satisfying $\|d_{n}\|\leq 2^{-n},$ and $\|F({{c^{2}_{n}}}d_{n})\|>\|G(c_{n})\|+n$ for all natural $n$. By assumptions, the limit $a_{0}=\sum_{n\geq 1}c_{n}d_{n}$ belongs to $\mathcal{A},$ $\|a_{0}\|\leq 1,$ and $a_{0}c_{m}=c_{m}a_{0}=c_{m}^{2}d_{m}$ for all $m\in\mathbb{N}$. Our hypotheses give
@@
Under some extra hypotheses on the bimodule, we conclude that the class of generalized Jordan derivations of the third type are in fact generalized derivation of first/second/third type.
@@
Let $\mathcal{A}$ be a C-algebra and let $\mathcal{M}$ be an essential Banach $\mathcal{A}$-bimodule. Then every generalized Jordan derivation of the third type $F:\mathcal{A}\rightarrow\mathcal{M}$ is a generalized derivation of the second type (and of course the first and third type), that is, there exist $\xi\in\mathcal{M}^{**}$ such that
@@
Furthermore, if $\mathcal{M}$ coincides with $\mathcal{A}$ or $\mathcal{A}^{*}$, then $\xi$ actually commutes all elements in $\mathcal{A}$.
@@
Since $\mathcal{M}$ is assumed to be essential, we can prove that $F$ is a generalized derivation of the second type by combining that $F$ is continuous (cf. Theorem 3.1) and Theorem 1.2. Namely, just observe that given $a,b\in\mathcal{A}_{sa}$ with $ab=0$, we can take $d\in\mathcal{A}$ with $d^{2}=a$ and $db=bd=0$. Then
$bF(a)b=bF(d^{2})b=b\left(2d\circ G(d)\right)b=0.$
@@
for all $a,b\in\mathcal{A}$ (compare the proof of Theorem 3.1). Then, in this case, $G$ is actually linear and continuous. Namely, we can apply similar ideas to those in the proof of Proposition 2.5, to see that the mappings of the form $x\mapsto(G(\cdot)\circ b)(x)=G(x)\circ b=F(x\circ b)-x\circ G(b)$ are linear and continuous on $\mathcal{A}$. Since for every $a,b\in\mathcal{A}$ with $\|b\|\leq 1$ we have $\|G(a)\circ b\|=\|F(a\circ b)-a\circ G(b)\|\leq\|F\|\|a\|+\|G(\cdot)\circ a\|,$ the family $\{G(\cdot)\circ b:\|b\|\leq 1\}$ is pointwise bounded, and hence uniformly bounded by the uniform boundedness principle. Thus there exists $K>0$ such that $\|G(\cdot)\circ b\|\leq K$ for all $\|b\|\leq 1$. Let $(u_{j})_{j}$ be a bounded approximate unit in $\mathcal{A}$. For each $a\in\mathcal{A}$, the net $F(a\circ u_{j})$ converges to $F(a)$ in norm by the continuity of $F$, and the net $G(a)\circ u_{j}$ tends to $G(a)$ in norm. Consequently $G$ is linear. We also know that $\|G(a)\|=\lim_{j}\|G(a)\circ u_{j}\|\leq K\|a\|,$ note that the first equality holds because $\mathcal{M}$ is essential (see, for example, [8, Theorem 1.15]). Thus, $G$ is continuous.
Having in mind that $F$ is a continuous generalized derivation on $\mathcal{A}$, and by the weak-density of $\mathcal{A}$ in $\mathcal{A}^{**}$ and the separate weak-continuity of the product of the latter von Neumann algebra, $F^{**}$ is a generalized derivation on $\mathcal{A}^{**},$ and thus there exists $\xi\in\mathcal{M}^{**}$ satisfying
@@
Let us prove the final statement. Assume first that $\mathcal{M}=\mathcal{A}$, then a similar argument, via weak-density of $\mathcal{A}$ in $\mathcal{A}^{**}$ and separate weak-continuity of the product in $\mathcal{A}^{**}$ and the identity in (3.1), shows that

### v2

@@ 2. $(ii)$
We shall next show that $I$ is a closed ideal of $\mathcal{A}$. Namely, take $a\in I$ and $b\in\mathcal{A}$. Clearly, the mapping $FL_{a}L_{b}:\mathcal{A}\rightarrow\mathcal{M}$, $c\mapsto F(abc)$ is continuous, and so $ab=ba\in I$. This means that $I$ is an ideal of $\mathcal{A}$ (recall that $\mathcal{A}$ is commutative). It is well-known that $I$ must be self-adjoint [31, Corollary 4.2.10].
Next we show that $I$ is norm-closed. Suppose that $(a_{n})_{n}$ is a sequence in $I$ converging to an element $a\in\mathcal{A}$. By the equality $I=J$, in order to show that $a\in I$, it suffices to show that the mapping $L_{a}H:\mathcal{A}\rightarrow\mathcal{M}$ is continuous. It follows from Lemma 2.3$(i)$ that the mapping $L_{c}H=:\mathcal{A}\rightarrow\mathcal{M}$ is linear for every $c\in\mathcal{A}$. Since $(a_{n})$ is a sequence in $I$, the linear mapping $L_{a_{n}}H:\mathcal{A}\rightarrow\mathcal{M}$ is continuous for all $n\in\mathbb{N}$. It is clear that $\lim_{n\rightarrow\infty}L_{a_{n}}H(c)=\lim_{n\rightarrow\infty}{a_{n}}H(c)=aH(c)=L_{a}H(c),$ for every $c\in\mathcal{A}$, and hence, by the uniform boundedness principle, we obtain that $L_{a}H$ is a continuous linear mapping, and so $a\in J=I$.
@@
Suppose, on the contrary, that ${\mathcal{A}}/{I}$ is an infinite-dimensional C-algebra. Another classical result in C-algebra theory (see [31, Exercise 4.6.13]) proves that in such a case there exists an infinite sequence $(b_{n}+I)_{n}$ of mutually orthogonal non-zero positive elements in ${\mathcal{A}}/{I}$, that is, $(b_{n}+I)(b_{k}+I)=0,$ for all $n\neq k$, $b_{n}\geq 0$ and $b_{n}\neq 0$ for all $n$. By [31, Exercise 4.6.20] we can always lift the sequence $(b_{n}+I)_{n}$ to a sequence $(c_{n})_{n}$ of mutually orthogonal non-zero positive elements in $\mathcal{A}$ satisfying $\pi(c_{n})=b_{n}+I$, for all $n$, where $\pi:\mathcal{A}\rightarrow{\mathcal{A}}/{I}$ is the canonical projection. It follows from the fact that $\pi$ is a -homomorphism that $\pi(c_{n}^{2})=\pi(c_{n})^{2}=(b_{n}+I)^{2}\neq 0$ (observe that $\|(b_{n}+I)^{2}\|=\|b_{n}+I\|^{2}$), and thus $c_{n}^{2}\notin I$ for all $n$. Therefore the mapping $FL_{{c^{2}_{n}}}:\mathcal{A}\rightarrow\mathcal{M}$ must be unbounded. By replacing $c_{n}$ with $\frac{c_{n}}{\|c_{n}\|}$, we can assume that $\|c_{n}\|=1$ for all $n$. By the unboundedness of the mapping $FL_{{c^{2}_{n}}}$, there exists $d_{n}\in\mathcal{A}$ satisfying $\|d_{n}\|\leq 2^{-n},$ and $\|F({{c^{2}_{n}}}d_{n})\|=\|FL_{{c^{2}_{n}}}(d_{n})\|>\|G(c_{n})\|+n$ for all natural $n$ –observe that the sequence $(\|G(c_{n})\|)_{n}$ does not produce any obstacle here–. The elements $c_{n}$ and $d_{n}$ have been chosen to guarantee that the series $\sum_{n\geq 1}c_{n}d_{n}$ is (absolutely) convergent, and its limit $a_{0}=\sum_{n=1}^{\infty}c_{n}d_{n}$ satisfies $\|a_{0}\|\leq 1$ and $c_{m}a_{0}=c_{m}^{2}d_{m}$ for all $m\in\mathbb{N}$. By the hypotheses on $(F,G,H)$ we deduce that
@@
Since $F|_{I}$ is continuous and $\mathcal{A}/I$ is finite-dimensional, we deduce from Lemma 2.6 that $F$ is continuous.
@@
The previous theorem generalizes the classical results by Ringrose [39, Theorem 2] and Sakai [44]. Let $F:\mathcal{A}\to\mathcal{M}$ be a linear mapping from a C-algebra to a Banach $\mathcal{A}$-bimodule. Suppose $F$ is a generalized derivation of the first type, that is, we can find a derivation $d:\mathcal{A}\to\mathcal{M}^{**}$ such that $F(ab)=F(a)b+ad(b)$ ($a,b\in\mathcal{A}$). Clearly, the mapping $F:\mathcal{A}\to\mathcal{M}\subseteq\mathcal{M}^{**}$ is a generalized derivation of the third type, and thus continuous by Theorem 2.8. Consequently every generalized derivation of the first type from a C-algebra to a Banach $\mathcal{A}$-bimodule is continuous.
The following corollary summarizes some of the conclusions obtained up to now.
@@ 1. $(a)$
    $F$ is a generalized derivation of the first type, i.e., there is a derivation $d:\mathcal{A}\to\mathcal{M}^{**}$ such that $F(ab)=F(a)b+ad(b)$ ($a,b\in\mathcal{A}$).
@@ 2. $(b)$
    $F$ is a generalized derivation of the second type, i.e., there is an element $\xi\in\mathcal{M}^{**}$ such that $F(ab)=F(a)b+aF(b)-a\xi b$ ($a,b\in\mathcal{A}$).
@@ 3. $(c)$
    $F$ is a generalized derivation of the third type, i.e., there are two (non-necessarily linear) maps $G,H:\mathcal{A}\to\mathcal{M}$ such that $F(ab)=G(a)b+aH(b)$ ($a,b\in\mathcal{A}$).
Furthermore, if any of the equivalent statements holds, the mapping $F$ is automatically continuous.
Let ${Y}$ and ${Z}$ be Banach spaces and let $T:{Y}\rightarrow{Z}$ be a linear mapping. The separating space of $T$ is the set
$\mathfrak{S}(T)=\Big{\{}z\in{Z}\ :\ \exists\ (y_{n})_{n}\subseteq{Y}\ such\ that\ (y_{n})_{n}\rightarrow 0,\hbox{ and }(T(y_{n}))_{n}\rightarrow z\Big{\}}.$
@@
Let $\mathcal{A}$ be a C-algebra, and let $\mathcal{M}$ be a Banach $\mathcal{A}$-bimodule. Suppose that $F,G,H:\mathcal{A}\rightarrow\mathcal{M}$ are three mappings, with $F$ linear, such that $(F,G,H)$ behaves like a derivation. The the following statements hold:
@@ 3. $(iii)$
    If $\mathcal{M}=\mathcal{A}^{*}$, then $F,G$ and $H$ are bounded linear maps.
@@
$(i)$ According to Theorem 2.8, $F$ is a continuous generalized derivation of the second type, and it follows from Lemma 2.3$(ii)$ that the mappings $G$ and $H$ are linear. We shall show that $G$ and $H$ are continuous. Let $m_{0}\in\mathfrak{S}(H)\subseteq\mathcal{M}$. Then, there exists a sequence $\{b_{n}\}\subseteq\mathcal{A}$ such that $\lim_{n\rightarrow\infty}b_{n}=0$ and $\lim_{n\rightarrow\infty}H(b_{n})=m_{0}$. For any arbitrary element $a\in\mathcal{A}$, we have
@@
Let $\mathcal{A}$ be an algebra and let $\mathcal{M}$ be an $\mathcal{A}$-bimodule. We recall that a linear mapping $F:\mathcal{A}\rightarrow\mathcal{M}$ is called an l-generalized derivation if there exists a mapping $F_{l}:\mathcal{A}\rightarrow\mathcal{M}$ such that $F(ab)=F(a)b+aF_{l}(b)$ for all $a,b\in\mathcal{A}$. Similarly, $F$ is called an r-generalized derivation if there exists a mapping $F_{r}:\mathcal{A}\rightarrow\mathcal{M}$ such that $F(ab)=F_{r}(a)b+aF(b)$ for all $a,b\in\mathcal{A}$. In [21, Example 2.6], we can find an example of an $r$-generalized derivation that is not an $l$-generalized derivation. Clearly, $l$- and $r$-generalized derivations are generalized derivations of the third type. The next result is therefore a straightforward consequence of Theorem 2.8 and Corollary 2.10.
@@ 2. $(ii)$
In [13] Freiberger and Mathieu considered an additional notion of “generalized derivation”. Let $\mathcal{A}$ be a unital C-subalgebra of the unital C-algebra $\mathcal{B}$. A linear map $\delta:\mathcal{A}\to\mathcal{B}$ is called a ternary-generalized derivation if, for all $x,y,z\in\mathcal{A}$ we have
$\delta(xyz)=\delta(xy)z-x\delta(y)z+x\delta(yz).$
These authors showed that $\delta$ is a ternary-generalized derivation if, and only if, there is a derivation $d:\mathcal{A}\to\mathcal{B}$ such that $\delta(xz)=d(x)z+x\delta(z)$ for all $x,z\in\mathcal{A}$. Obviously, ternary-generalized derivations are generalized derivations of the second and third type, and so our results can be also applied in this case.
@@ 3. Automatic continuity of generalized Jordan derivation of the third type
Appropriate types of generalized Jordan derivations whose domain is a C-algebra are studied in this section. We shall mainly focus on the automatic continuity and the relationships between the different types. Contrary to the conclusions in the previous section, the natural definition for generalized Jordan derivations of the third type defines a strict subclass of the set of generalized Jordan derivations of the second type (cf. Proposition 3.2 and Remark 3.3). Generalized Jordan derivations of the second type have been considered in [6, 7, 26]. Recall that a linear mapping $F$ from a Banach algebra $\mathcal{A}$ to a Banach $\mathcal{A}$-bimodule $\mathcal{M}$ is a generalized Jordan derivation of the second type if there exists $\xi\in\mathcal{M}^{**}$ such that
@@
where $a\circ b=\frac{1}{2}(ab+ba)$ and $U_{a,b}(\xi)=\frac{1}{2}(a\xi b+b\xi a)$. If $\xi=0,$ we find the usual notion of Jordan derivation. Classic results by Ringrose and Johnson assure that every Jordan derivation from a C-algebra to a Banach bimodule is a derivation.
There are some other attempts to define generalized Jordan derivations in the literature, According to [5], an additive (in this note we shall assume linearity) mapping $T$ from a Banach algebra $\mathcal{A}$ into a left $A$-module $M$ is called a Jordan left derivation if $T(a^{2})=2aT(a)$ for every $a\in\mathcal{A}$. It is shown in [5] that the existence of non-zero Jordan left derivations from a prime ring $\mathcal{R}$ into a 6-torsion free left $\mathcal{R}$-module implies that $\mathcal{R}$ is a commutative ring. Vukman introduced in [45] the notion of $(m,n)$-Jordan derivation.
Jing and Lu coined for the first time the term generalized Jordan derivations in [28]. We shall say that a linear mapping $F$ from a Banach algebra $\mathcal{A}$ to a Banach $\mathcal{A}$-bimodule $\mathcal{M}$ is a generalized Jordan derivation in the sense of Jing and Lu if there exists a (linear) Jordan derivation $\tau:\mathcal{A}\to\mathcal{M}$ satisfying
@@
In this definition the Jordan structure is only considered on the domain algebra. It is worthwhile to mention that if $\mathcal{A}$ and $\mathcal{M}$ are both unital, by taking $b=\mathbf{1}$ in the previous identity we get $F(a)=F(\mathbf{1})a+\tau(a)$ ($a\in\mathcal{A}$), and if $\tau$ is, in fact, a derivation (as in the case of C-algebras), $F$ is a generalized derivation of the first type.
@@
The following theorem establishes the automatic continuity of every generalized Jordan derivation of the third type when the domain is a C-algebra.
@@
We are now in a position to apply Lemma 2.7 to guarantee that $F|_{I}:{I}\to\mathcal{M}$ is continuous. The continuity of $F$ will follow from Lemma 2.6 if we show that ${\mathcal{A}}/{I}$ is finite-dimensional. Otherwise, as in the proof of Theorem 2.8, there exists an infinite sequence $(b_{n}+I)_{n}$ of mutually orthogonal non-zero positive elements in ${\mathcal{A}}/{I}$, that is, $(b_{n}+I)(b_{k}+I)=0,$ for all $n\neq k$, $b_{n}\geq 0$ and $b_{n}\neq 0$ for all $n$ (see [31, Exercise 4.6.13]). Choose, a sequence $(c_{n})_{n}$ of mutually orthogonal norm-one positive elements in $\mathcal{A}$ satisfying $c_{n}+I=b_{n}+I$, $\|c_{n}^{2}+I\|=\|(b_{n}+I)^{2}\|=\|b_{n}+I\|^{2}\neq 0$, for all $n$ (cf. [31, Exercise 4.6.20]). Clearly, the element $c_{n}^{2}$ does not belong to $I$, and hence the mapping $x\mapsto F(c_{n}^{2}x)$ is unbounded. We can therefore pick $d_{n}\in\mathcal{A}$ satisfying $\|d_{n}\|\leq 2^{-n},$ and $\|F({{c^{2}_{n}}}d_{n})\|>\|G(c_{n})\|+n$ for all natural $n$. By assumptions, the limit $a_{0}=\sum_{n\geq 1}c_{n}d_{n}$ belongs to $\mathcal{A},$ $\|a_{0}\|\leq 1,$ and $a_{0}c_{m}=c_{m}a_{0}=c_{m}^{2}d_{m}$ for all $m\in\mathbb{N}$. Our hypotheses give
@@
Under some extra hypotheses on the bimodule, we can conclude that every generalized Jordan derivation of the third type on a C-algebra is, in fact, a generalized derivation of the first, second, and third type.
@@
Let $\mathcal{A}$ be a C-algebra and let $\mathcal{M}$ be an essential Banach $\mathcal{A}$-bimodule. Then every generalized Jordan derivation of the third type $F:\mathcal{A}\rightarrow\mathcal{M}$ is a generalized derivation of the second type (and, of course, of the first and third type), that is, there exist $\xi\in\mathcal{M}^{**}$ such that
@@
Furthermore, if $\mathcal{M}$ coincides with $\mathcal{A}$ or $\mathcal{A}^{*}$, the element $\xi$ can be chosen satisfying $\xi a=a\xi$, for all $a\in\mathcal{A}$.
@@
for all $a,b\in\mathcal{A}$ (compare the proof of Theorem 3.1).
Since $\mathcal{M}$ is assumed to be essential, we can prove that $F$ is a generalized derivation of the second type by combining that $F$ is continuous (cf. Theorem 3.1) and Theorem 1.2. To see this, just observe that given $a,b\in\mathcal{A}_{sa}$ with $ab=0$, we can take $d\in\mathcal{A}$ with $d^{2}=a$ and $db=bd=0$. Therefore,
$bF(a)b=bF(d^{2})b=b\left(2d\circ G(d)\right)b=0,$
which concludes the proof of the first statement.
We shall next show that $G$ is actually linear and continuous. Namely, we can apply similar ideas to those in the proof of Proposition 2.5, to see that the mappings of the form $x\mapsto(G(\cdot)\circ b)(x)=G(x)\circ b=F(x\circ b)-x\circ G(b)$ are linear and continuous on $\mathcal{A}$. Since for every $a,b\in\mathcal{A}$ with $\|b\|\leq 1$ we have $\|G(a)\circ b\|=\|F(a\circ b)-a\circ G(b)\|\leq\|F\|\|a\|+\|G(\cdot)\circ a\|,$ the family $\{G(\cdot)\circ b:\|b\|\leq 1\}$ is pointwise bounded, and hence uniformly bounded by the uniform boundedness principle. Thus, there exists $K>0$ such that $\|G(\cdot)\circ b\|\leq K$ for all $\|b\|\leq 1$. Let $(u_{j})_{j}$ be a bounded approximate unit in $\mathcal{A}$. For each $a\in\mathcal{A}$, the net $F(a\circ u_{j})$ converges to $F(a)$ in norm by the continuity of $F$, and the net $G(a)\circ u_{j}$ tends to $G(a)$ in norm. Consequently $G$ is linear. We also know that $\|G(a)\|=\lim_{j}\|G(a)\circ u_{j}\|\leq K\|a\|,$ note that the first equality holds because $\mathcal{M}$ is essential (see, for example, [20, Theorem 32.22] and [8, Theorem in page 108, § 1.15]). Thus, $G$ is continuous.
Having in mind that $F$ is a continuous generalized derivation of the second type on $\mathcal{A}$, Theorem 2.11 in [2] assures that $F^{**}$ is a generalized derivation of the second type on $\mathcal{A}^{**},$ and thus there exists ${\xi}\in\mathcal{M}^{**}$ satisfying ${\xi}=\mathbf{1}{\xi}\mathbf{1},$ and
@@
Observe that, a priori, the element ${\xi}$ belongs to $\mathcal{M}^{****}$, however, since $\mathbf{1}F^{**}(\mathbf{1})\mathbf{1}=\mathbf{1}\widehat{\xi}\mathbf{1}=\widehat{\xi}$ lies in $\mathcal{M}^{**}$.
Let us prove the final statement. Assume first that $\mathcal{M}=\mathcal{A}$. The weak-density of $\mathcal{A}$ in $\mathcal{A}^{**}$ and the separate weak-continuity of the product in $\mathcal{A}^{**}$ combined with the identity in (3.1) prove that